OpenStudy (anonymous):

whats the answer to a=1/2 bh?

OpenStudy (anonymous):

what is the area of a triangle

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

No I mean a = 1/2 bh IS the area of a triangle. im not sure what your question is

OpenStudy (anonymous):

thats all it says

OpenStudy (kymber):

Are there instructions prior to this? Are there choices? We definitely need more information.

OpenStudy (anonymous):

u got a diagram attached to this q?

OpenStudy (anonymous):

Suppose you are trying to find the height of a triangle given the area and the length of the base. Rewrite the area formula below to solve for the height of the triangle

OpenStudy (anonymous):

thanks

OpenStudy (kymber):

Oh, okay. So you want to get 'h' alone. 'h' is the variable that's the height. How can you make it so it's h=...?

OpenStudy (anonymous):

distibute idk

OpenStudy (anonymous):

i wasnt supposed to get into 8th grade i did just barely

OpenStudy (anonymous):

this is what you have so far...\[area = \frac{base \times height}{2}\]

OpenStudy (anonymous):

HINT:- your first move should be to do with the 2

OpenStudy (anonymous):

put them both over 2?

OpenStudy (anonymous):

@kymber carry on :)

OpenStudy (anonymous):

i really have no clue what to do im retarded

OpenStudy (anonymous):

you are not retarded

OpenStudy (kymber):